OpenStudy (anonymous):

x + y = 15 4x + 6y = 72

OpenStudy (anonymous):

I need help again finding x then plugging it in to find y

OpenStudy (kc_kennylau):

Use Wolfram|Alpha

OpenStudy (kc_kennylau):

(1)*6: 6x+6y=90 ---(3) (3)-(2): 2x=18 --> x=9

OpenStudy (anonymous):

subtract y from both sides in the first equation and then substitute for x.

OpenStudy (kc_kennylau):

method of @AlexisEyvonne also works :)

OpenStudy (anonymous):

I need help with another please :)

OpenStudy (anonymous):

3x + 2y = 475 x + 2y = 180

OpenStudy (kc_kennylau):

(1)-(2): 2x=295 --> x=147.5